At a maximum level of loudness, the power output of a 75-piece orchestra radiated as sound is 72.2 W. What is the intensity of these sound waves to a listener who is sitting 22.3 m from the orchestra? Answer in units of W/m^2
Answer #1

Intensity = Power / area = 72.2 / [4Ï€ R^2]

4Ï€ R^2 = 4*22*22.3*22.3 /7 = 6249.13 m^2

I = 72.2 / 6249.13


I =11.55 * 10 ^ -3 W /m^2
Answer #2

intensity = power /area

intensity = 72.2 / (pi * 22.3^2)

intensity = 0.04621 W / m^2